Grade 9Mechanics

A water drop falls at regular intervals from tap 5m above ground. The third drop is leaving the tap at the instant the first job touches the ground how far above the ground is the second drop at that instant

surendra surendra
7 Years agoGrade 9

1 Answer

Arun
ApprovedApproved Tutor Answer7 Years ago
Dear Surendra
 
 

Height of tap = 5m and (g) = 10 m/sec2.

For the first drop, 5 = ut +  gt 2

= (0 × t) +  × 10 t2 v= 5tor t= 1 or t = 1.

It means that the third drop leaves after one second of the first drop. Or, each drop leaves after every 0.5 sec.

Distance covered by the second drop in 0.5 sec

 

        = (0.5)2 = 1.25m

        Therefore, distance of the second drop above the ground

= 5 - 1.25 = 3.75 m.

 

Regards

Arun